# Federated Learning as a Service for Hierarchical Edge Networks with Heterogeneous Models

### v1

@@ 1 Introduction
Federated learning (FL) [22] is a distributed learning algorithm designed for privacy-aware applications. Without compromising the user’s privacy, FL aggregates trained local models to build a new global model. In each training round, all client devices receive the global model from the cloud server, train it using local data, and send the updated model back to the server for building a new global model in the aggregation process (Figure 1). Compared with traditional centralized training, FL alleviates concerns about direct data violation [12], as the client’s original data is never transmitted to the server in the training process. In recent years, Machine Learning as a Service (MLaaS) has gained widespread attention, proved by the increasing demand for cloud-based machine learning platforms and large-scale analytics across various industries [28]. However, the concerns of widely collecting privacy data have also grown [7]. FL, as a decentralized machine learning, has the capability to face the challenges of privacy concerns. In 2020, Nicolas et al. proposed Federated Learning as a Service (FLaaS) for permission and privacy management [13]. In recent years, FL has been applied in various applications such as healthcare systems [2], prediction maintenance for Internet of Things (IoT) devices [6], and edge networks [24, 21].
@@
Data heterogeneity among client devices is One of the major challenges in FL [18]. In practical FL scenarios, the client’s data is highly dependent on user behavior. As a result, training data on clients’ devices may have significantly different distributions [9]. This phenomenon is called non-Independent and Identical (non-IID) distributed data, which has a significant impact on the global model’s performance [20]. Device heterogeneity in the computational hierarchy is another existing problem of FL. Most proposed FL frameworks require all participants to use the same model architecture regardless of their computational resources. However, in practical FL scenarios such as Google Gboard [31], mobile clients could have various hardware configurations ([16], [14], and [29]). In this setting, running the same model on all client devices is challenging, if it is not impossible. Devices with low computational resources are incapable of finishing the training process on time [14], which increases the average time for a training round as the server can perform the model aggregation when it receives trained models from all devices.
@@
To address these research problems, various methods have been proposed to aggregate heterogeneous models within conventional Federated Learning (FL) systems [23], [5]. However, most of these methods cannot reach the target performance when encountering non-IID data [19]. In this paper, we introduce an aggregation method for heterogeneous models in three-level hierarchical FL frameworks (IoT/edge/cloud) called HAF-Edge. In this computational hierarchy, IoT devices are clustered based on some specific requirements (data/hardware configurations). Each cluster is connected to an edge server that receives the trained models and aggregates them to create an edge-aggregated model. All edge devices are connected to a cloud server that receives edge-aggregated models with different structures. The could server performs the second level of aggregation by using knowledge sharing and based on similarity in structures of different models to build a global model for each edge server (Figure  1). HAF-Edge not only leverages the advantages of hierarchical federated learning, potentially reducing the need for the Internet to aggregate trained IoT models at the edge [25], but also addresses the challenge of aggregating edge models with different settings at the cloud level. By connecting a cluster of IoT nodes with the same setting to an edge server using local communication technologies (e.g., Bluetooth or Zigbee), the need for using the Internet in creating edge-aggregated models is eliminated [25] (reduces communication cost). HAF-Edge addresses the challenge of aggregating heterogeneous models at the cloud by adopting a strategy called Max-Common [27] in hierarchical edge networks. At the cloud level, we extract parts of the knowledge from different model settings to enhance the global model’s aggregation process. Our experimental results demonstrate that models trained with IID data tend to have a greater distance from the global model compared to scenarios where non-IID data is used. Based on this observation, we propose a distance-based weighting aggregation approach and apply it to the aggregation process at the edge level to improve performance for applications with non-IID data. The main contributions of our work are summarized as follows:
@@
  • 2) A distance-based weighting aggregation approach is proposed, which makes HAF-Edge enables to achieve a better performance facing non-IID data compared to the state-of-art FL frameworks.
@@ 2 Related Works
FL [22] is a decentralized model training method that was proposed for training neural networks in a distributed fashion without sharing the user’s data. However, the existence of non-IID data among client devices in practical FL scenarios is one of the most important problems that reduces the performance of the global model [34]. In vanilla FedAvg, only data volume is considered in the aggregation process. As a result, in practical FL scenarios with non-IID data, regardless of data distribution, the same weighting coefficient is considered for trained models of two client devices that had data with the same size but with completely different labels. Broadly aggregating the models trained with low-quality data can have a serious impact on the performance of the global model in FL [4]. Some proposed methods start to shift the focus to the quality of client data or models. FAIR proposed in [4] improves the performance of the global model by filtering the local models with low quality. Moreover, a game-theoretic FL framework is proposed in [33] to activate clients to provide the local models trained with high-quality data.
In practical IoT FL scenarios, the expected local model architecture may differ among client devices due to heterogeneous hardware configurations, tasks, and personal demands [11] [18] [32]. To address these challenges, methods such as FedMD [17] and MHAT [10] have been proposed based on knowledge distillation [8]. However, in these methods, a public dataset is created, which can compromise client data privacy. In contrast to knowledge distillation, MaxCommon [27] aggregates common layers of heterogeneous models to share learned features among participants. This approach targets heterogeneous model architectures and refines the aggregation of local updates to specific layers rather than entire models. By aggregating as many common layers as possible from different clients’ models, MaxCommon maximizes knowledge sharing. The method employs data quantity-based aggregation to merge common layers from various clients’ updates, ensuring that a client’s model with a larger training dataset contributes more to the aggregated common layers. However, this approach does not consider the distribution of local training data.
@@ 3.1 Distance-based Weighting
  • Configuration 1: In total three participants, two participants have balanced distributed data while the other has skewed distributed data, data volume are the same across these devices.
  • Configuration 2: In total three participants, one participant has balanced distributed data while the rest have skewed distributed data, data volume are the same across these devices.
@@
$d(p,q)=\sqrt{(p_{1}-q_{1})^{2}+(p_{2}-q_{2})^{2}+(p_{3}-q_{3})^{2}+\cdots+(p_{n}-q_{n})^{2}}$ (1)
where $d(p,q)$ is the Euclidean distance between $p$ and $q$, where $p$ and $q$ are matrices in n dimensions. FedAvg is utilized to aggregate three local models from client devices in both Configuration 1 and Configuration 2. The distance between the local models of three client devices and the aggregated global model from the previous round is calculated using the Euclidean distance metric.
@@ 3.2 Edge Aggregation
$E_{i}^{t}$ An edge aggregated model from edge server $i$ at round $t$ (the aggregated model from edge server $i$ contains $i$ layers in total)
@@
$E_{i}^{1}=\sum_{k\in K_{i}}\frac{N_{k}}{\sum_{k\in K_{i}}N_{k}}L_{k}^{1}$ (3)
where $N_{k}$ is the total number of data used to train the local model in the client $k$. After receiving edge-aggregated models, the cloud server creates new global models with different architectures by aggregating common layers from various edge-aggregated models. New global models are distributed to matching edge servers. The edge server propagates the received new global model to connected client devices. Hence, starting from the second round, the distance-based weighting aggregation is applied at the edge server. The edge aggregated model $E_{i}^{t}$ at round $t(t>1)$ from edge server $i$ is:
$E_{i}^{t}=\sum_{k\in K_{i}}\frac{d(L_{k}^{t},G_{i}^{t-1})}{\sum_{k\in K_{i}}d(L_{k}^{t},G_{i}^{t-1})}L_{k}^{t}$ (4)
@@
if $t=1$ then
     $E_{i}^{t}=\sum_{k\in K_{i}}\dfrac{N_{k}}{\sum_{k\in K_{i}}N_{k}}L_{k}^{t}$
else
     $E_{i}^{t}=\sum_{k\in K_{i}}\dfrac{d(L_{k}^{t},G_{i}^{t-1})}{\sum_{k\in K_{i}}d(L_{k}^{t},G_{i}^{t-1})}L_{k}^{t}$
end if
return $E_{i}^{t}$
@@ 3.3 Cloud Aggregation
$G_{i}^{t}=Gl_{1}^{t}\oplus Gl_{2}^{t}\oplus Gl_{3}^{t}\oplus\cdots\oplus Gl_{i}^{t}$ (7)
@@
Algorithm 2 The Aggregation at Cloud Server. $n$ is the number of layers contained by the edge server with the highest layer count. $X_{i}$ is a set of edge servers whose aggregated models contain layer $i$. $E_{x}^{t}$ is an edge aggregated model from edge server $x$ at round $t$. $l_{i}^{x}$ is layer $i$ from $E_{x}^{t}$. $Gl_{i}^{t}$ is the global layer $i$. $K_{x}$ is the set of client devices connected to edge server $x$. $N_{k}$ is the total number of data used for training in the client $k$. $G_{i}^{t}$ is the global model for edge server $i$. $t$ is the round counter.
for $i=1,2,3,\dots,n$ do
     for $x\in X_{i}$ do
         $l_{i}^{x}=Extract(E_{x},i)$
     end for
     $Gl_{i}^{t}=\sum_{x\in X_{i}}\dfrac{\sum_{k\in K_{x}}N_{k}}{\sum_{x\in X_{i}}\sum_{k\in K_{x}}N_{k}}l_{i}^{x}$
end for
for $i=1,2,3,\dots,n$ do
     $G_{i}^{t}=Gl_{1}^{t}\oplus Gl_{2}^{t}\oplus Gl_{3}^{t}\oplus\cdots\oplus Gl_{i}^{t}$
end for
return $G_{1}^{t},G_{2}^{t},G_{3}^{t},\dots,G_{n}^{t}$
@@ 4 Implementation and Evaluation
In our study, we evaluate the effectiveness of HAF-Edge using MNIST and FMNIST by comparing its performance with two baselines: vanilla FedAvg [22] and MaxCommon strategy [27]. The TensorFlow library of Python is used for implementation. In our experiment, we use heterogeneous models with different numbers of dense layers. Table.2 represents the foundation model structure. The x-nn represents the model which has x number of dense layers except the output layer. As an example, 1nn and 5nn represent models that have 1 dense layer and 5 dense layers, respectively. To verify the effectiveness of our method, three different scenarios are applied to evaluate the performance of HAF-Edge. Table.3 shows the detailed settings of three experiment scenarios. 1NN and 3NN models are used in Scenario 1 and Scenario 2. In Scenario 3, 5 different model architectures from 1NN to 5NN are employed. To generate non-IID training sets for the client devices in three scenarios, MNIST and FMNIST datasets are sorted according to labels and partitioned into several subsets based on the required number of non-IID client devices in different scenarios, and each subset is regarded as one non-IID client. The client device with IID data contains all labels in three scenarios.
@@ 4.2 Evaluation Results
For Scenario 3, the number of edge servers and model architecture types is increased compared to Scenario 1 and Scenario 2. Figures 7 and 8 present the test accuracy trends for global 1NN and 5NN models. MNIST dataset is used in Scenario 3. The results of HAF-Edge are better in both test accuracy and convergence speed compared to the FedAvg and MaxCommon strategies. In all cases, the models trained by HAF-Edeg can achieve at least 10% accuracy higher than models trained by FedAvg and MaxCommon. For the 5NN model, the maximum accuracy of HAF-Edge is about 80%, while the best performance of both FedAvg and MaxCommon reach to 60% accuracy.
@@ 5 Conclusion
Non-IID data and model heterogeneity are two important challenges of FLaaS. In this paper, we present HAF-Edge, a three-level hierarchical federated learning where client devices with the same model settings are clustered and connected to an edge server that aims to solve these problems. In HAF-Edge, MaxCommon strategy is adopted at the cloud server to extract and aggregate common layers from models with different architectures. A distance-based weighting aggregation is proposed and applied at edge servers to reduce the impact of data non-IIDness, which is based on our observation that the Euclidean distance between local models trained with IID data and the global model is higher compared with the local models trained with non-IID data. We also evaluate the performance of HAF-Edge based on MNIST, FMNIST datasets under multiple scenarios. The results indicate that HAF-Edge outperforms FedAvg and MaxCommon Strategy under scenarios with non-IID data and heterogeneous models.

### v2

@@ 1 Introduction
Federated learning (FL) [22] is a distributed learning algorithm designed for privacy-aware applications. Without compromising the user’s privacy, FL aggregates trained local models to build a new global model [24]. In each training round, all client devices receive the global model from the cloud server, train it using local data, and send the updated model back to the server for building a new global model in the aggregation process (Figure 1). Compared with traditional centralized training, FL alleviates concerns about direct data violation [12], as the client’s original data is never transmitted to the server in the training process. In recent years, Machine Learning as a Service (MLaaS) has gained widespread attention, proved by the increasing demand for cloud-based machine learning platforms and large-scale analytics across various industries [29]. However, the concerns of widely collecting privacy data have also grown [8]. FL, as a decentralized machine learning, has the capability to face the challenges of privacy concerns. In 2020, Nicolas et al. proposed Federated Learning as a Service (FLaaS) for permission and privacy management [13]. In recent years, FL has been applied in various applications such as healthcare systems [2], prediction maintenance for Internet of Things (IoT) devices [7], and edge networks [25, 21].
@@
Data heterogeneity among client devices is one of the major challenges in FL [17]. In practical FL scenarios, the client’s data is highly dependent on user behavior. As a result, training data on clients’ devices may have significantly different distributions [10]. This phenomenon is called non-Independent and Identical (non-IID) distributed data, which has a significant impact on the global model’s performance [19, 4]. Device heterogeneity in the computational hierarchy is another existing problem of FL. Most proposed FL frameworks require all participants to use the same model architecture regardless of their computational resources. However, in practical FL scenarios such as Google Gboard [31], mobile clients could have various hardware configurations [14]. In this setting, running the same model on all client devices is challenging, if it is not impossible. Devices with low computational resources are incapable of finishing the training process on time [14], which increases the average time for a training round as the server can perform the model aggregation when it receives trained models from all devices.
@@
To address these research problems, various methods have been proposed to aggregate heterogeneous models within conventional Federated Learning (FL) systems [23]. However, most of these methods cannot reach the target performance when encountering non-IID data [18]. In this paper, we introduce an aggregation method for heterogeneous models in three-level hierarchical FL frameworks (IoT/edge/cloud) called HAF-Edge. In this computational hierarchy, IoT devices are clustered based on some specific requirements (data/hardware configurations), such as the facilities in a smart city or industry of IoT. Each cluster is connected to an edge server that receives the trained models and aggregates them to create an edge-aggregated model. All edge servers are connected to a cloud server that receives edge-aggregated models with different structures. The cloud server performs the second level of aggregation by using knowledge sharing and based on similarity in structures of different models to build a global model for each edge server. HAF-Edge not only leverages the advantages of hierarchical federated learning, potentially reducing the need for the Internet to aggregate trained IoT models at the edge [26], but also addresses the challenge of aggregating edge models with different settings at the cloud level. By connecting a cluster of IoT nodes with the same setting to an edge server using local communication technologies (e.g., Bluetooth or Zigbee), the need for using the Internet in creating edge-aggregated models is eliminated [26] (reduces communication cost). HAF-Edge addresses the challenge of aggregating heterogeneous models at the cloud by adopting a strategy called MaxCommon [28] in hierarchical edge networks. At the cloud level, we extract parts of the knowledge from different model settings to enhance the global model’s aggregation process. Our experimental results demonstrate that models trained with IID data tend to have a greater distance from the global model compared to scenarios where non-IID data is used. Based on this observation, we propose a distance-based weighting aggregation approach and apply it to the aggregation process at the edge level to improve performance for applications with non-IID data. The main contributions of our work are summarized as follows:
@@
  • 2) A distance-based weighting aggregation approach is proposed, which makes HAF-Edge enables to achieve a better performance facing non-IID data compared to the state-of-the-art FL frameworks.
@@ 2 Related Works
FL [22] is a decentralized model training method that was proposed for training neural networks in a distributed fashion without sharing the user’s data. However, the existence of non-IID data among client devices in practical FL scenarios is one of the most important problems that reduces the performance of the global model [33]. In vanilla FedAvg, only data volume is considered in the aggregation process. As a result, in practical FL scenarios with non-IID data, regardless of data distribution, the same weighting coefficient is considered for trained models of two client devices that had data with the same size but with completely different labels. Broadly aggregating the models trained with low-quality data can have a serious impact on the performance of the global model in FL [6]. Some proposed methods start to shift the focus to the quality of client data or models. FAIR proposed in [6] improves the performance of the global model by filtering the local models with low quality.
In practical IoT FL scenarios, the expected local model architecture may differ among client devices due to heterogeneous hardware configurations, tasks, and personal demands [11, 17, 32]. To address these challenges, FedMD [16] has been proposed based on knowledge distillation [9]. However, in the methods using knowledge distillation, a public dataset is created, which can compromise client data privacy. In contrast to knowledge distillation, some methods aggregate common layers of heterogeneous models to share learned features among participants, such as MaxCommon [28] and Rank-Based Lora Aggregation [5]. MaxCommon targets heterogeneous model architectures and refines the aggregation of local updates to specific layers rather than entire models. By aggregating as many common layers as possible from different clients’ models, MaxCommon maximizes knowledge sharing. The method employs data quantity-based aggregation to merge common layers from various clients’ updates, ensuring that a client’s model with a larger training dataset contributes more to the aggregated common layers. However, this approach does not consider the distribution of local training data.
@@ 3.1 Distance-based Weighting
  • Configuration 1: In total three participants, one participant has balanced distributed data while the rest have skewed distributed data, data volume are the same across these devices.
  • Configuration 2: In total three participants, two participants have balanced distributed data while the other has skewed distributed data, data volume are the same across these devices.
@@
$d(p,q)=\sqrt{(p_{1}-q_{1})^{2}+(p_{2}-q_{2})^{2}+(p_{3}-q_{3})^{2}+\cdots+(p_{n}-q_{n})^{2}},$ (1)
where $d(p,q)$ is the Euclidean distance between $p$ and $q$, $p$ and $q$ are matrices in n dimensions. FedAvg is utilized to aggregate three local models from client devices in both Configuration 1 and Configuration 2. The distance between the local models of three client devices and the aggregated global model from the previous round is calculated using the Euclidean distance metric.
@@ 3.2 Edge Aggregation
$E_{i}^{t}$ An edge-aggregated model from edge server $i$ at round $t$ (the aggregated model from edge server $i$ contains $i$ layers in total)
@@
$E_{i}^{1}=\sum_{k\in K_{i}}\frac{N_{k}}{\sum_{k\in K_{i}}N_{k}}L_{k}^{1},$ (3)
where $N_{k}$ is the total number of data used to train the local model in the client $k$. After receiving edge-aggregated models, the cloud server creates new global models with different architectures by aggregating common layers from various edge-aggregated models. New global models are distributed to matching edge servers. The edge server propagates the received new global model to connected client devices. Hence, starting from the second round, the distance-based weighting aggregation is applied at the edge server. The edge-aggregated model $E_{i}^{t}$ at round $t(t>1)$ from edge server $i$ is:
$E_{i}^{t}=\sum_{k\in K_{i}}\frac{d(L_{k}^{t},G_{i}^{t-1})}{\sum_{k\in K_{i}}d(L_{k}^{t},G_{i}^{t-1})}L_{k}^{t}.$ (4)
@@
1:if $t=1$ then
2:     $E_{i}^{t}=\sum_{k\in K_{i}}\dfrac{N_{k}}{\sum_{k\in K_{i}}N_{k}}L_{k}^{t}$
3:else
4:     $E_{i}^{t}=\sum_{k\in K_{i}}\dfrac{d(L_{k}^{t},G_{i}^{t-1})}{\sum_{k\in K_{i}}d(L_{k}^{t},G_{i}^{t-1})}L_{k}^{t}$
5:end if
6:return $E_{i}^{t}$
@@ 3.3 Cloud Aggregation
$G_{i}^{t}=Gl_{1}^{t}\oplus Gl_{2}^{t}\oplus Gl_{3}^{t}\oplus\cdots\oplus Gl_{i}^{t},$ (7)
@@
Algorithm 2 The Aggregation at Cloud Server. $n$ is the number of layers contained by the edge server with the highest layer count. $X_{i}$ is a set of edge servers whose aggregated models contain layer $i$. $E_{x}^{t}$ is an edge-aggregated model from edge server $x$ at round $t$. $l_{i}^{x}$ is layer $i$ from $E_{x}^{t}$. $Gl_{i}^{t}$ is the global layer $i$. $K_{x}$ is the set of client devices connected to edge server $x$. $N_{k}$ is the total number of data used for training in the client $k$. $G_{i}^{t}$ is the global model for edge server $i$. $t$ is the round counter.
1:for $i=1,2,3,\dots,n$ do
2:     for $x\in X_{i}$ do
3:         $l_{i}^{x}=Extract(E_{x},i)$
4:     end for
5:     $Gl_{i}^{t}=\sum_{x\in X_{i}}\dfrac{\sum_{k\in K_{x}}N_{k}}{\sum_{x\in X_{i}}\sum_{k\in K_{x}}N_{k}}l_{i}^{x}$
6:end for
7:for $i=1,2,3,\dots,n$ do
8:     $G_{i}^{t}=Gl_{1}^{t}\oplus Gl_{2}^{t}\oplus Gl_{3}^{t}\oplus\cdots\oplus Gl_{i}^{t}$
9:end for
10:return $G_{1}^{t},G_{2}^{t},G_{3}^{t},\dots,G_{n}^{t}$
@@ 4 Implementation and Evaluation
In our study, we evaluate the effectiveness of HAF-Edge using MNIST and FMNIST which are recognized as benchmarks in FL systems for image classification applications by comparing its performance with two baselines: vanilla FedAvg [22] and MaxCommon strategy [28]. The TensorFlow library of Python is used for implementation. In our experiment, we use heterogeneous models with different numbers of dense layers. Table.2 represents the foundation model structure. The x-nn represents the model which has x number of dense layers except the output layer. As an example, 1nn and 5nn represent models that have 1 dense layer and 5 dense layers, respectively. To verify the effectiveness of our method, three different scenarios are applied to evaluate the performance of HAF-Edge. Table.3 shows the detailed settings of three experiment scenarios. 1NN and 3NN models are used in Scenario 1 and Scenario 2. In Scenario 3, 5 different model architectures from 1NN to 5NN are employed. To generate non-IID training sets for the client devices in three scenarios, MNIST and FMNIST datasets are sorted according to labels and partitioned into several subsets based on the required number of non-IID client devices in different scenarios, and each subset is regarded as one non-IID client. The client device with IID data contains all labels in three scenarios.
@@ 4.2 Evaluation Results
For Scenario 3, the number of edge servers and model architecture types is increased compared to Scenario 1 and Scenario 2. Figures 7 and 8 present the test accuracy trends for 5 different global models (1NN, 2NN, 3NN, 4NN, 5NN). MNIST dataset is used in Scenario 3. The results of HAF-Edge are better in both test accuracy and convergence speed compared to the FedAvg and MaxCommon strategies. In all cases, the models trained by HAF-Edge can achieve at least 10% accuracy higher than models trained by FedAvg and MaxCommon. For the 5NN model, the maximum accuracy of HAF-Edge is about 80%, while the best performance of both FedAvg and MaxCommon reach to 60% accuracy.
@@ 5 Conclusion
Non-IID data and model heterogeneity are two important challenges of FLaaS. In this paper, we present HAF-Edge, a three-level hierarchical federated learning where client devices with the same model settings are clustered and connected to an edge server that aims to solve these problems. In HAF-Edge, MaxCommon strategy is adopted at the cloud server to extract and aggregate common layers from models with different architectures. A distance-based weighting aggregation is proposed and applied at edge servers to reduce the impact of data non-IIDness, which is based on our observation that the Euclidean distance between local models trained with IID data and the global model is larger compared with the local models trained with non-IID data. We also evaluate the performance of HAF-Edge based on MNIST, FMNIST datasets under multiple scenarios. The results indicate that HAF-Edge outperforms FedAvg and MaxCommon Strategy under scenarios with non-IID data and heterogeneous models.